I am having problems with this and a few other word problems, so here is this first on meny. I'll see if I have done wright. It go as follows'
The total length of two boards is 94 5/8". The difference between there two length's is 24 1/4". Find the length's of the two board's.
Good Luck..& thanks

Answer by stanbon(75887)

The total length of two boards is 94 5/8". The difference between there two length's is 24 1/4". Find the length's of the two board's.
---------------
Let the shorter board length be "x".
Then the longer board length is "x+24 1/4"
--------------------
Equation:
sharter + longer = total
x + x+ 24 1/4 = 94 5/8
--------------------
Adjust the fraction:
2x + 24 2/8 = 94 5/8
2x = 70 3/8
x = 35 3/16 (length of the shorter board)
-------
x + 24 1/4 = 35 3/16 + 24 4/16 = 59 7/16 (length of the longer board)
================================================
Cheers,
Stan H.